Consider a homogeneous-product Cournot oligopoly with four firms. Suppose that the inverse demand function is P(Q) = 64 – Q. Suppose that firms incur a constant marginal cost c = 4. Characterize the equilibrium of the game in which all firms simultaneously choose quantity.

Suppose that firms 1 and 2 consider merging and that there are synergies leading to marginal costs cm < c. Characterize the new market equilibrium. At what level of cm are the two firms indifferent whether to merge or not?

Does such a merger in which the two firms are just indifferent between merging and not merging increase consumer surplus? Briefly explain your response.
Answer #1

8) Q = EU = 4 +92 +93 +94 MC = 4, Now mi= (P-MC) q i = (P-4) U c un = ( 64 -4 - Yi- Q-1) vi = (60-Q-i) qi - q2 Qi- sag Brie a/ Now Date BRM. anmaq (64-cm-93-94)-29m =0 q (64-cm - 93-94)/2 Now T3 = (P-4) 93 = (64-4-qm-93-94) 93 AN = (60-9M-94 )q 3 - 9Date TM= [(42-30)/17 7/2 each firm of Merged from gets = MTOM (42-3cm ) 32. P=64-48 &= 12x4 = 48 I when No Memer, - 16. - 8 T